OpenStudy (tia_sweetheart17):

square root of 80p^3. show work

OpenStudy (anonymous):

\[\sqrt{80p^3}\]= \[\sqrt{16} \sqrt{p^2}\] times \[\sqrt{5} \sqrt{p}\] = \[4p \sqrt{5p}\]